OpenStudy (anonymous):

a flagpole casts a shadow 28 feet long. a person standing nearby casts a shadow eight feet long. if a person is six feet tall,how tall is the flagpole

OpenStudy (31356):

21

OpenStudy (31356):

Medal?

OpenStudy (31356):

Set up a proportion

OpenStudy (anonymous):

yelpp shes correct you will have 28/8=x/6 then you will have 8x=168 divide by 8 then x=21 :) good job 31356

OpenStudy (31356):

Thanks!

OpenStudy (31356):

Thanks for the medal! :)

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

your welcome